OpenStudy (anonymous):

Solve the equation (y-3)(y-5)=0

OpenStudy (bahrom7893):

3 and 5

OpenStudy (anonymous):

3 and 5

OpenStudy (bahrom7893):

come on it's already factored

OpenStudy (bahrom7893):

3-3 = 0, making the whole thing 0. Same for 5-5

OpenStudy (texaschic101):

y - 3 = 0 (add 3 to both sides) y = 3 y - 5 = 0 (add 5 to both sides) y = 5